OpenStudy (anonymous):

Solve. -14 + x² = 5x

OpenStudy (anonymous):

x will hv two values x=7 x=-2

OpenStudy (anonymous):

x=-2,7

OpenStudy (anonymous):

Consider the equivalent equation : \[x ^{2}-5x-14=0\] introduce \[\Delta=b^2-4ac\] (with a=2, b=5, c =-14). Solutions are \[(-b \pm \sqrt{\Delta})/2a\]

OpenStudy (anonymous):

b=-5, sorry.

OpenStudy (anonymous):

it can simply be done using middle term splitting..